A producer of new corn seed guarantees that 85% of the seeds will germinate. A gardener plants 10 of the seeds. What is the probability that;

a) None will grow

b) Six will grow

c) At least nine will grow


1
Expert's answer
2022-04-29T14:40:05-0400

Let "X=" the number of corn seed which growed: "X\sim Bin (n, p)."

Given "p=0.85, q=1-p=0.15, n=10."

a)


"P(X=0)=\dbinom{10}{0}(0.85)^0(0.15)^{10-0}"

"\approx5.7665\times10^{-9}"

b)


"P(X=6)=\dbinom{10}{6}(0.85)^6(0.15)^{10-6}"

"\approx0.0401"

c)


"P(X\ge9)=P(X=9)+P(X=10)"

"=\dbinom{10}{9}(0.85)^9(0.15)^{10-9}+\dbinom{10}{10}(0.85)^{10}(0.15)^{10-10}"

"\approx0.5443"